40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 26 17:50:22 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement 
    Reply  Wed Jan 27 18:19:45 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement 
       Reply  Thu Jan 28 00:29:30 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement PLL_OLG.pdfFreq_noise_comparison.pdf
          Reply  Mon Feb 1 18:11:25 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement Free_running_frequency_noise_comparison.pdf
       Reply  Fri Jan 29 14:43:47 2016, Steve, Update, Green Locking, Innolight laser is 10 years old inno1W.pdf
          Reply  Fri Jan 29 16:55:27 2016, rana, Update, Green Locking, Innolight laser is 10 years old 
             Reply  Sat Jan 30 00:12:22 2016, gautam, Update, Green Locking, Innolight laser is 10 years old RIN_comparison.pdf
                Reply  Sat Jan 30 09:56:24 2016, Koji, Update, Green Locking, Innolight laser is 10 years old 
                   Reply  Mon Feb 1 15:16:28 2016, gautam, Update, Green Locking, Innolight laser is 10 years old 
                      Reply  Mon Feb 1 15:43:18 2016, Koji, Update, Green Locking, Innolight laser is 10 years old 
                         Reply  Fri Feb 5 15:02:13 2016, gautam, Update, Green Locking, X-end NE cable 
          Reply  Fri Feb 12 11:10:49 2016, Steve, Update, Green Locking, Innolight laser is 10 years old inno1W.jpg
          Reply  Mon Mar 21 14:29:32 2016, Steve, Update, Calibration-Repair, 1W Innolight laser repair diagnoses 
             Reply  Thu Mar 24 07:56:09 2016, Steve, Update, Calibration-Repair, NO Noise Eater for 1W Innolight 
                Reply  Mon Apr 11 17:03:41 2016, Steve, Update, Calibration-Repair, 1W Innolight repair completed 
Message ID: 11969     Entry time: Mon Feb 1 18:11:25 2016     In reply to: 11956
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: Lightwave frequency noise measurement 

Before distrubing the beat setup with the spare Lightwave laser, I wanted to see if I could resolve the apparent difference in behaviour between the measured free running noise of the spare Lightwave laser and my earlier measurements with the existing X and Y end lasers above ~5kHz. So I redid the measurement, but this time, on Eric's suggestion, while taking spectra on the SR785, I was careful to maintain the same "CH1 input range" while measuring the control signal spectrum and the measurement noise spectra. The level used was -20dBvpk. I think the measured spectrum shape now makes sense - above ~4kHz, the SR560 noise means that the SNR is poor and so we can only trust the spectra up to this value (the spectra for the end lasers are from earlier measurements where I did not take care to keep the input range constant). Anyways, I think the conclusion is that the spare Lightwave seems to have a free-running frequency noise that is approximately a factor of 3 worse than the Lightwave laser at the Y-end, though this may be because I didn't take the measurement at the optimal operating conditions (diode current, power etc). But I guess this is tolerable and that we can go ahead with the planned swapping out of the existing Innolight at the X-end with this laser. 

I will now move the Lightwave laser off the PSL table onto the SP table where I will do some beam characterization and see if I can come up with a satisfactory mode-matching solution for the swap. I've borrowed a beam profiler from the TCN lab for this purpose.

Attachment 1: Free_running_frequency_noise_comparison.pdf  103 kB  | Show | Show all
ELOG V3.1.3-